The measure of an angles supplement is 76 less than the measure of the angle. Find the measure of the angle and its supplement.
Answer by cleomenius(959)

2x + 76 = 180
2x = 104
x = 52 degrees.
52 + 76 = 128 degrees.
52 + 128 = 180 degrees.
This does check.
Celomenius.